OpenStudy (secret-ninja):

Medals! Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18

OpenStudy (secret-ninja):

A. 12x + 4y = -12 15x = -30 B. 8x + 4y = -4 14x = -10 C. 6x - 8y = -36 -6y = -42 D. 6x - y = -15 3y = 9